\begin{thebibliography}{99}
%\setlength{\biblabelsep}{2mm}
%{\setlength{\baselineskip}{0.75\baselineskip}
\bibitem{Almgren1966}
F. J. Almgren, Jr. Some interior regularity theorems for minimal surfaces and an extension of Berenstein's theorem, Ann. of Math. {\bf{84}} (1966), 277-292.
\bibitem{Antonelli-Xu}
G. Antonelli and K. Xu, New spectral Bishop-Gromov and Bonnet-Myers theorems and applications to isoperimetry, to appear in J. Eur. Math. Soc., available at arXiv:2405.08918v2(2024).
\bibitem{Bel23}
C. Bellettini, Extensions of Schoen-Simon-Yau and Schoen-Yau theorems via iteration \`{a} la De Giorgi, Invent. Math. {\bf{240}} (2025), no. 1, 1–34.
\bibitem{Bernstein1927}
S. Bernstein, \"{U}ber ein geometrisches Theorem und seine Anwendung auf die partiellen Differentialgleichungen vom elliptischen Typus, Math. Z. {\bf{26}} (1927), 551-558.
\bibitem{BDE}
E. Bombieri, E. De Giorgi and E. Giusti, Minimal cones and the Bernstein problem, Invent. Math. {\bf{7}} (1969), 243-268.
\bibitem{Bray97}
H. L. Bray, The Penrose inequality in general relativity and volume comparison theorems involving scalar curvature, Proquest LLC, Ann Arbor, MI, 1997, Thsis(Ph.D.)-Stanford University.
\bibitem{Brendle}
S. Brendle, The isoperimetric inequality for a minimal submanifold in Euclidean space. J. Amer. Math. Soc. {\bf{34}} (2021), 595-603.
\bibitem{Cao-Shen-Zhu1997}
H. D. Cao, Y. Shen and S. H. Zhu, The structure of stable minimal hypersurfaces in $\RR^{n+1}$, Math. Res. Lett. {\bf{4}} (1997), 637-644.
\bibitem{Catino-Mastrolia-Roncoroni}
G. Catino, P. Mastrolia and A. Roncoroni, Two rigidity results for stable minimal hypersurfaces, Geom. Funct. Anal. {\bf{34}} (2024), 1-18. %https://doi.org/10.1007/s00039-024-00662-1.
\bibitem{general soap bubbles}
O. Chodosh and C. Li, Generalized soap bubbles and the topology of manifolds with positive scalar curvature, Ann. of Math. {\bf{199}} (2024), no. 2, 707-740.
\bibitem{Chodosh-Li}
O. Chodosh and C. Li, Stable minimal hypersurfaces in $\bR^4$, Acta Math. {\bf{233}} (2024), no. 1, 1-31.
\bibitem{Chodosh-Li-anisotropic}
O. Chodosh and C. Li, Stable anisotropic minimal hypersurfaces in $\mathbb{R}^4$, Forum Math. Pi {\bf{11}} (2023), Paper NO. e3, 22pp.
\bibitem{Chdosh-Li-Stryker-positive cueved}
O. Chodosh, C. Li and D. Stryker, Complete stable minimal hypersurfaces in positively curved 4-manifolds, J. Eur. Math. Soc. {\bf{28}} (2026), no. 4, 1457–1487.
\bibitem{Chodosh-Li-Minter-Stryker-5bernstein}
O. Chodosh, C. Li, P. Minter and D. Stryker, Stable minimal hypersurfaces in $\bR^{5}$, to appear in Ann. of Math., available at arXiv:2401.01492v3(2024).
\bibitem{Colding-Minicozzi02}
T. H. Colding and W. P. Minicozzi $\secondfund$, Estimates for parametric elliptic integrands, Int. Math. Res. Not. {\bf{2002}} (2002), no. 6, 291-297.
\bibitem{De Giorgi1965}
E. De Giorgi. Una estensione del teorema di Berenstein, Ann. Scuola Norm. Sup. Pisa Cl. Sci. {\bf{19}} (1965), 79-85.
\bibitem{do Carmo-Peng1979}
M. do Carmo and C. K. Peng, Stable complete minimal surfaces in $\bR^3$ are planes, Bull. Am. Math. Soc. (N.S.) {\bf{1}} (1979), no. 6, 903-906.
\bibitem{FCS80}
D. Fischer-Colbrie and R. Schoen, The structure of complete stable minimal surfaces in 3-manifolds of nonnegative scalar curvature, Commun. Pure Appl. Math. {\bf{33}} (1980), 199-211.
\bibitem{Fleming1962}
W. H. Fleming, On the oriented Plateau problem, Rend. Circ. Mat. Palermo (2) {\bf{11}} (1962), 69-90.
\bibitem{GL}
R. Gulliver and H. B. Lawson, Jr., The structure of stable minimal hypersurfaces near a singularity, in Geometric measure theory and the calculus of variations(Arcata, Calif., 1984), Proceedings of Symposia in Pure Mathematics vol. 44(American Mathematical Society, Providence, RI, 1986), 213-237.
\bibitem{Gromov2019Fourlecture}
M. Gromov, Positive curvature, macroscopic dimension, spectral gaps and higher signatures.
In Functional analysis on the eve of the 21st century, Vol. II (New Brunswick, NJ, 1993), volume 132 of Progr. Math., pages 1–213. Birkh\"auser Boston, Boston, MA, 1996.
\bibitem{Hong-cmc nonexis}
H. Hong, Z. Yan, Rigidity of CMC hypersurfaces in 5-and 6-manifolds, arXiv:2405.06867v3(2024).

\bibitem{HLW24}
H. Hong, H. Li and G. Wang, On $\delta$-Stable Minimal Hypersurfaces in $\mathbb{R}^{n+1}$,
arXiv:2407.03222(2024).

\bibitem{Jenkins1961}
H. Jenkins, On $2$-dimensional variational problems in parametric form, Arch. Ration. Mech. Anal. {\bf{8}} (1961), 181-206.
\bibitem{F.H-Lin}
F. H. Lin, Extimates for surfaces which are stationary for an elliptic parameter integral, J. Partial Differential Equations {\bf{3}} (1990), no. 3, 78-92.
\bibitem{Mazet-6Bernstein}
L. Mazet, Stable minimal hypersurfaces in $\bR^6$, arXiv:2405.14676v1(2014).
\bibitem{Sobolev-ineq}
J. H. Michael and L. M. Simon, Sobolev and mean-value inequalities on generalized sumanifolds of $\bR^{n}$, Comm. Pure. Appl. Math. {\bf{26}} (1973), 361-379.
\bibitem{Mooney}
C. Mooney, Entire solutions to equations of minimal surfaces type in six dimensions, J. Eur. Math. Soc. {\bf{24}} (2022), no. 12, 4353-4361.
\bibitem{Mooney-Yang}
C. Mooney, Y. Yang, The anisotropic Bernstein problem, Invent. Math. {\bf{235}} (2023), 211-232.
\bibitem{Pogorelov1981}
A. V. Pogorelov, On the stability of minimal surfaces, Sov. Math., Dokl. {\bf{24}} (1981), 274-276.
\bibitem{Schoen-Simon-Yau}
R. Schoen, L. Simon, and S. T. Yau, Curvature estimates for minimal hypersurfaces, Acta Mathematica. {\bf{134}} (1975), no. 3-4, 275-288.
\bibitem{Shen-Ye96}
Y. Shen and R. Ye, On stable minimal surfaces in manifolds of positive bi-Ricci curvatures, Duke Math. J. {\bf{85}} (1996), no. 1, 109-116.
%\bibitem{Shen-Ye-diam}
%Y. Shen and R. Ye, On the geometry and topology of manifolds of positive bi-Ricci curvature, arXiv e-prints(1997), arXiv: 9708014.
\bibitem{Sim68}
J. Simons, Minimal varieties in Riemannian manifolds, Ann. of Math. {\bf{88}} (1968), no. 22, 62-105.
\bibitem{Simon1977}
L. Simon, On some extensions of Bernstein's theorem, Math. Z. {\bf{154}} (1977), 265-273.
\bibitem{White91}
B. White, Existence of smooth embedded surfaces of prescribed genus that minimize parametric even elliptic functionals on 3-manifolds, J. Differential Geom. {\bf{33}} (1991), no. 2, 413-443.
\bibitem{Winklmann}
S. Winklmann, Pointwise curvature estimates for stable $F$-minimal hypersurfaces, Ann. Inst. H. Poincar\'{e} C Anal. Non Lin\'{e}aire {\bf{22}} (2005), no. 5, 543-555.
\bibitem{zhujintian2021}
J. Zhu, Width estimate and doubly warped product, Trans. Amer. Math. Soc. {\bf{374}} (2021), no. 2, 1497-1511.

\end{thebibliography}

\documentclass[11pt]{amsart}

\usepackage{graphicx}
\usepackage{mathrsfs}
\usepackage{color}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{mathrsfs}
\usepackage[margin = 1.4in] {geometry}
\usepackage[hyperindex,breaklinks]{hyperref}
\usepackage{xcolor}

\newtheorem{prop}{Proposition}
\newtheorem{theo}[prop]{Theorem}
\newtheorem*{theo*}{Theorem}
\newtheorem{lemm}[prop]{Lemma}
\newtheorem{coro}[prop]{Corollary}

\newtheorem{rema}[prop]{Remark}
\newtheorem*{Ackn}{Acknowledgements}

\theoremstyle{definition}

\makeatletter

\newcommand{\Rmnum}[1]{\expandafter\@slowromancap\romannumeral #1@}
\makeatother

\newcommand{\RR}{\mathbf{R}}

\newcommand{\cA}{\mathcal A}

\newcommand{\bR}{\mathbb{R}}
\newcommand{\bS}{\mathbb{S}}

\DeclareMathOperator{\dis}{dist}

\DeclareMathOperator{\tr}{tr}

\DeclareMathOperator{\Vol}{Vol}

\DeclareMathOperator{\Ric}{Ric}
\DeclareMathOperator{\biRic}{biRic}

\DeclareMathOperator{\Div}{div}

\DeclareMathOperator{\secondfund}{II}
\DeclareMathOperator{\Rm}{Rm}

\newcommand{\bangle}[1]{\left\langle #1 \right\rangle}

\title[stable anisotropic minimal hypersurfaces]{Stable anisotropic minimal hypersurfaces \\ in $\bR^{5}$ and $\bR^{6}$}
\thanks{C.X. is supported by NSFC (Grant No. 12271449, 12126102) and the Natural Science Foundation of Fujian Province of China (Grant No. 2024J011008).}

\author{Jia Li}
\address{School of Mathematical Sciences\\
		Xiamen University\\
		361005, Xiamen, P.R. China}
\email{lijiamath@stu.xmu.edu.cn}
\thanks{Key words and phrases: Anisotropic minimal hypersurface, stable  minimal hypersurface, Bernstein's problem, $\mu$-bubble.}
\author{Chao Xia}
\address{School of Mathematical Sciences\\
		Xiamen University\\
		361005, Xiamen, P.R. China}
\email{chaoxia@xmu.edu.cn}

\begin{document}

\begin{abstract}
In this paper, we prove that a complete, two-sided, stable anisotropic minimal immersed  hypersurface in $\bR^{5}$ or $\bR^{6}$ is flat, provided the anisotropic area functional is $C^4$-close to the area functional. 
\end{abstract}
\maketitle

\section{Introduction}
The classical Bernstein problem asks whether a minimal graph in $\bR^{n+1}$ is flat. The picture is nowadays complete that it has been answered affirmatively for $n\le 7$ by Bernstein \cite{Bernstein1927}, Fleming \cite{Fleming1962}, Almgren \cite{Almgren1966}, de Giorgi \cite{De Giorgi1965} and Simons \cite{Sim68} and negatively for $n\ge 8$ by Bombieri-de Giorgi-Giusti \cite{BDE}. The stable Bernstein problem asks whether a complete {two-sided} stable minimal hypersurface in $\bR^{n+1}$ is flat. It has been solved for $n=2$ by do Carmo-Peng and {Fischer Colbrie-Schoen} and Pogorelov, in \cite{do Carmo-Peng1979,FCS80,Pogorelov1981} respectively, and for $n\le 6$ by Schoen-Simon-Yau \cite{Schoen-Simon-Yau} under the Euclidean volume growth assumption (see also Bellettini \cite{Bel23}). 
Quite recently, there are important progress for the stable Bernstein problem. Precisely, it has been first solved by Chodosh-Li \cite{Chodosh-Li} for $n=3$, and two different proofs have been given by Chodosh-Li \cite{Chodosh-Li-anisotropic} and Catino-Mastrolia-Roncoroni \cite{Catino-Mastrolia-Roncoroni}. In particular,  Chodosh-Li's proof \cite{Chodosh-Li-anisotropic}, which is based on $\mu$-bubble method, yields the intrinsic Euclidean volume growth for a complete stable minimal hypersurface. This method has been successfully applied to solve stable Bernstein problem in $\bR^5$ by Chodosh-Li-Minter-Stryker \cite{Chodosh-Li-Minter-Stryker-5bernstein} and in $\bR^6$ by Mazet \cite{Mazet-6Bernstein}. In view of Bombieri-de Giorgi-Giusti's counterexample in $\bR^8$. The only remaining unsolved case is stable Bernstein problem in $\bR^7$.

For a two-sided immersion $M^n\to \bR^{n+1}$ with chosen unit normal vector field $\nu$, we consider the anisotropic area functional or parametric elliptic integral
$$\mathcal{A}_F(M):=\int_{M}F(\nu)d\mu,$$
where $F:\bR^{n+1}\setminus\{0\}\to(0,\infty)$ is a positive $C^3$ function which is $1$-homogeneous, i.e. $F(tz)=t F(z)$ for $t>0$,
and is elliptic which means $D^2F(z)$ is a positive definite endomorphism for all $z\in\mathbb{S}^n$. We say $M$ is $F$-minimal if $\frac{d}{dt}|_{t=0}\mathcal{A}_F(M_{t})=0$ for all compactly supported variations of $M$  and that $M$ is stable if in addition $\frac{d^2}{dt^2}|_{t=0}\mathcal{A}_F(M_{t})\geq 0$ for such variations. In particular, when $F(\nu)=|\nu|$, $\mathcal{A}_F$ reduces to the classical area functional, and the $F$-minimal and the stable $F$-minimal  hypersurfaces are classical minimal and stable minimal hypersurfaces, respectively. 
The $\mathcal{A}_F$ functionals have recently attracted significant attention
due to their practical applications and theoretical significance.

The (stable) anisotropic Bernstein problem asks the similar questions about flatness of $F$-minimal graphs or stable $F$-minimal hypersurfaces. The anisotropic Bernstein problem for graph cases has been resolved in the affirmative by Jenkins \cite{Jenkins1961} in $\bR^3$ and Simon \cite{Simon1977} in $\bR^4$. The stable anisotropic Bernstein problem has been resolved in the affirmative in $\bR^3$ assuming quadratic area growth by White \cite{White91}.
However, there are non-flat $\mathcal{A}_F$-minimizers in $\bR^{n+1}, n\geq 3$ by Mooney \cite{Mooney} and Mooney-Yang \cite{Mooney-Yang}. 
On the other hand, under the assumption of sufficient closeness of $\mathcal{A}_F$ to area, the stable anisotropic Bernstein problem has been resolved in the affirmative in $\bR^3$ by Lin \cite{F.H-Lin} (see also Colding-Minicozzi \cite{Colding-Minicozzi02}) and in $\bR^4$ by Chodosh-Li \cite{Chodosh-Li-anisotropic}, as well as in $\bR^{n+1}, 2\le n\le 5$ by Winklmann \cite{Winklmann} under the additional assumption of Euclidean volume growth following the seminal work of Schoen-Simon-Yau \cite{Schoen-Simon-Yau}. 
In particular, Chodosh-Li \cite{Chodosh-Li-anisotropic} used the $\mu$-bubble method to get the Euclidean volume growth for stable $F$-minimal hypersurfaces in $\bR^4$. We would like to mention that the $\mu$-bubble method has been initiated by Gromov \cite{Gromov2019Fourlecture} for problems involving scalar curvature and has many applications.
\begin{theo}[\cite{Chodosh-Li-anisotropic}]
Assume that $F$ satisfies
  \begin{align*}
 |\xi|^2\leq D^2F(z)(\xi,\xi)\leq \sqrt{2}|\xi|^2,
 \end{align*}
 for all $\xi\in z^{\perp}$. 
 Let $M^3\to\bR^{4}$ be a complete, two-sided, simply-connected, stable $F$-minimal immersion. Then there is some constant $C(F)$ such that $${\rm Vol}(B_R(p))\leq C(F)R^3,$$ where  $B_R(p)$ denotes the geodesic ball of radius $R$ centered at $p\in M$.\par
\end{theo}

Recall that the stable Bernstein problem in $\bR^{n+1}, n=4, 5$ has been solved by Chodosh-Li-Minter-Stryker \cite{Chodosh-Li-Minter-Stryker-5bernstein} and Mazet \cite{Mazet-6Bernstein}, again using Chodosh-Li's \cite{general soap bubbles} $\mu$-bubble method and some other new ingredients. It is natural to ask whether Chodosh-Li's result in the anisotropic case is true for $\bR^{n+1}, n=4,5$. This is the main aim for this paper. 
\begin{theo}\label{main-theo}
   Let $n=4,5$ and assume that $F$ satisfies
  \begin{align}\label{cond-1}
 |\xi|^2\leq D^2F(z)(\xi,\xi)\leq (1+\epsilon_{n})|\xi|^2,
 \end{align}
 for all $\xi\in z^{\perp}$, with $\epsilon_{4}=\frac{3}{20}$ and $\epsilon_{5}=\frac{1}{1000}$. 
 Let $M^n\to\bR^{n+1}$ be a complete, two-sided, stable $F$-minimal immersion.
  Then there is some constant $C(F)$ such that $${\rm Vol}(B_R(p))\leq C(F)R^n.$$
\end{theo}
Combining with Winklmann's result \cite{Winklmann}, we get
\begin{coro}
   Let $n=4,5$ and assume that $\cA_F$ is $C^4$-sufficiently close to the area. Then any complete,  two-sided,  stable $F$-minimal immersed hypersurface in $\bR^{n+1}$ is flat.
\end{coro}
Here, we say that  $\cA_{F}$ is $C^{4}$-sufficiently close to the area functional if $$\|F-1\|_{C^{4}(\bS^{n})}\leq \sigma_n$$ for some fixed $\sigma_n>0$. 

Let us first illustrate the $\mu$-bubble method to the Euclidean volume growth for stable minimal hypersurfaces developed by Chodosh-Li \cite{Chodosh-Li-anisotropic} in $\bR^4$, by Chodosh-Li-Minter-Stryker \cite{Chodosh-Li-Minter-Stryker-5bernstein} in $\bR^5$ and by Mazet \cite{Mazet-6Bernstein} in $\bR^6$. The first step is to consider a conformal metric $\tilde{g}=r^{-2}g$, where $r$ is the Euclidean distance to $p\in M$, inspired by Gulliver-Lawson \cite{GL}. Then the stability of the minimal hypersurface can be transferred to the property of a positive spectral lower bound for the scalar curvature ($n=3$), or the bi-Ricci curvature ($n=4$) or  some weighted bi-Ricci curvature ($n=5$), for the induced metric $\tilde{g}$ on $M\setminus\{p\}$. The next step is to find a good $\mu$-bubble that satisfies a  positive spectral lower bound for Ricci curvature. This leads to a spectral Bishop-Gromov volume estimate by Chodosh-Li-Minter-Stryker \cite{Chodosh-Li-Minter-Stryker-5bernstein} in three dimensions and Antonelli-Xu \cite{Antonelli-Xu} in any dimensions, by using isoperimetric profile method originally due to Bray \cite{Bray97}. The last step is to convert the volume estimate of $\mu$-bubble to the original metric $g$ and the growth of volume of $M$ is controlled by isoperimetric inequality of Michael-Simon \cite{Sobolev-ineq} and Brendle \cite{Brendle}.

As mentioned above, Chodosh-Li's strategy for $\bR^4$ works for 
 the anisotropic setting, when $F$ satisfies the pinching condition (\ref{cond-1}). We see that the above strategy also works for $\bR^5$ and $\bR^6$ when $D^2F$ is enough pinched. 
The main observation is that $H^2\le \delta_{n}^2 |A|^2$ for some small $\delta_{n}$ when $D^2F$ is enough pinched, though the classical mean curvature $H$ does not vanish in the anisotropic case. We first prove that $M$ has only one end, generalizing Cao-Shen-Zhu's result \cite{Cao-Shen-Zhu1997}. Then similar to the classical case, we show that the stability of $F$-minimal hypersurface can be transferred to the property of a positive spectral lower bound for the bi-Ricci curvature ($n=4$) or  some weighted bi-Ricci curvature ($n=5$), for the induced conformal metric $\tilde{g}$. The other part of proof follows the line as above.

\ 

\section{PRELIMINARIES}

We recall the first and the second variational formulas for the anisotropic area functional, see for example \cite[Section A]{Chodosh-Li-anisotropic}.
Let $M^{n}\to\bR^{n+1}$ be a two-sided immersed hypersurface and $M_t$ be a compact supported variation fixing $\partial M$ with a variational vector field $f\nu$, $f\in  C^{1}_{0}(M\setminus \partial M)$. Then 
$$\frac{d}{dt}\big|_{t=0}\mathcal{A}_F(M_{t})=\int_{M}\Div_{M}(DF(\nu))f.$$
The $F$-mean curvature is defined by $$H_{F}:=\Div_{M}(DF(\nu))=tr_M(\Psi(\nu)S_M),$$ 
where $S_{M}$ is the shape operator of $M$, $D$ is the connection in $\bR^{n+1}$, and the map $\Psi(\nu):T\bR^{n+1}\to T\bR^{n+1}$ is defined by $\Psi(\nu):V\mapsto D^{2}F(\nu)[V,\cdot]$.
Thus $M$ is $F$-minimal if and only if $H_F=0$.
Let $M^{n}\to\bR^{n+1}$ be an $F$-minimal two-sided immersed hypersurface, denoting $\nabla$ by the induced connection on $M$, then
  $$\frac{d^2}{dt^2}\big|_{t=0}\cA_{F}(M_{t})=\int_{M}\bangle{\nabla f,\Psi(\nu)\nabla f}-\tr_{M}(\Psi(\nu)S_{M}^2)f^2.$$
Thus, when $F$ satisfies the pinching condition \eqref{cond-1} and $M$ is stable $F$-minimal, then 
\begin{eqnarray}\label{delta-stable}\int_{M}|\nabla f|^2-\frac{1}{1+\epsilon_{n}}|A|^2f^2\geq 0,
\end{eqnarray}
 for any $f\in C^{1}_{0}(M\setminus \partial M)$, where $A$ is the second fundamental form of the immersion. 
\begin{rema}
    We mention that $M$ which satisfying inequality like \eqref{delta-stable} is called $\delta$-stable in the literature. For recent developments on  $\delta$-stable minimal hypersurfaces, we refer to \cite{HLW24} and the reference therein.
\end{rema}

The first variational formula leads to the following isoperimetric inequality.
 \begin{prop}[\cite{Chodosh-Li-anisotropic}]\label{iso-perimetric}
    Let $M^{n}\to\bR^{n+1}$ be a $F$-minimal immersion and the image of $\partial M$ is connected in $B^{\bR^{n+1}}_R$ for some $R>0$. Then $$|M|\leq\frac{R\|F\|_{C^{1}(\bS^{n})}}{n\cdot\min\limits_{\nu\in\bS^{n}}F(\nu)}|\partial M|,$$
    where $\|F\|_{C^{m}(\bS^n)}:=\big(\sum\limits_{j=0}^{m}\|D^{(j)}F\|^2_{C^{0}(\bS^n)}\big)^{\frac{1}{2}}$, and $B^{\bR^{n+1}}_R$ denotes the standard  ball of radius $R$ in $\bR^{n+1}$.
 \end{prop}
The stability of $M$ and the Michael-Simon-Sobolev inequality together lead to the $L^2$-Sobolev inequality on $M$, which implies that $M$ has infinite volume.
\begin{prop}[\cite{Chodosh-Li-anisotropic}]\label{infinite volume}
  Let $M^{n}\to\bR^{n+1}$ be a two-sided complete, stable $F$-minimal hypersurface, and $K$ is a compact subset of $M^{n}$. Then each unbounded component of $M\setminus K$ has infinite volume.
 \end{prop}
 On the other hand, combining the $L^2$-Sobolev inequality and Proposition \ref{infinite volume}, the same arguments as used by Cao-Shen-Zhu  \cite{Cao-Shen-Zhu1997} imply the following.
 \begin{prop}[\cite{Chodosh-Li-anisotropic}]\label{least-two ends}
   Let $M^{n}\to\bR^{n+1}$ be a complete two-sided, stable $F$-minimal immersed hypersurface with at least two ends, then there is a bounded non-constant harmonic function $u$ on $M$ with finite Dirichlet energy.
 \end{prop}
 
The following lemma establishes a control of the mean curvature $H$ by the second fundamental form $A$ of $M$, under the assumption (\ref{cond-1}) that $D^{2}F(\nu)$ is sufficiently pinched.
\begin{lemm}\label{mean-estimate}
  Assume $F$ satisfies (\ref{cond-1}). Let $M^{n}\to \bR^{n+1}$ be an $F$-minimal immersion. Then  \begin{eqnarray}\label{Hsmall}H^2\leq \delta_{n}^2|A|^2,\end{eqnarray}
  where the constant $\delta_{n}^2=\frac{(n-1)\epsilon_{n}^2}{(1+\epsilon_{n})^2}$.
\end{lemm}

\begin{proof}
 If we diagonalise the second fundamental form $A$ at a given point $p$, and denote $\kappa_{i}$ and $e_i$ for the principal curvatures and the corresponding principal directions, then the $F$-mean curvature at $p$ can be written as $$H_F=\sum\limits_{i=1}^{n}a_{i}\kappa_{i},$$
  where $a_{i}=D^2F(\nu)[e_i, e_i]$. Without loss of generality, we assume that $a_{1}\leq a_{2}\leq \cdots\leq a_{n}$. The pinching assumption (\ref{cond-1}) implies $$1\leq a_{1}\leq a_{2}\leq \cdots\leq a_{n}\leq 1+\epsilon_{n}.$$
  Since $H_F=0$, we have $$\kappa_{n}=-\frac{a_{1}\kappa_{1}+a_{2}\kappa_{2}+\cdots+a_{n-1}\kappa_{n-1}}{a_{n}}.$$ Write $\beta_{1}=\frac{a_{1}}{a_{n}}$, $\beta_{2}=\frac{a_{2}}{a_{n}}$,$\cdots$, $\beta_{n-1}=\frac{a_{n-1}}{a_{n}}$, with $\frac{1}{1+\epsilon_{n}}\leq\beta_{1}\leq\beta_{2}\leq\cdots\leq\beta_{n}\leq 1$. Then, we obtain
$$H^2=(\kappa_{1}+\kappa_{2}+\cdots+\kappa_{n})^2=((1-\beta_{1})\kappa_{1}+(1-\beta_{2})\kappa_{2}+\cdots+(1-\beta_{n-1})\kappa_{n-1})^2.$$
Using Cauchy-Schwarz's inequality, we get
\begin{align*}
  \frac{H^2}{|A|^2}&=\frac{((1-\beta_{1})\kappa_{1}+(1-\beta_{2})\kappa_{2}+\cdots+(1-\beta_{n-1})\kappa_{n-1})^2}{\kappa_{1}^2+\kappa_{2}^2+\cdots+\kappa_{n-1}^{2}+\kappa_{n}^2}\\
&\leq\frac{(n-1)((1-\beta_{1})^2\kappa_{1}^2+(1-\beta_{2})^2\kappa_{2}^2+\cdots+(1-\beta_{n-1})^2\kappa_{n-1}^2)}{\kappa_{1}^2+\kappa_{2}^2+\cdots+\kappa_{n-1}^2}\\
&\leq (n-1)(1-\frac{1}{1+\epsilon_{n}})^2=\delta_{n}^2.
\end{align*}
\end{proof}

\section{one-endness of stable $F$-minimal hypersurfaces in $\bR^{n+1}$}
In this section, we prove that any $n$-dimensional complete two-sided stable $F$-minimal immersed hypersurface $M^{n}\to \bR^{n+1}$ must have exactly one end, if $D^{2}F$ is sufficiently pinched.
\begin{lemm}\label{one-end-improve}
Assume $F$ satisfies (\ref{cond-1}). Let $M^n$ be a complete two-sided, stable $F$-minimal immersion in $\bR^{n+1}$ and $u$ be a harmonic function on $M$. Then
\begin{align*}
\Lambda_{n}\int_{M}\varphi^2|A|^2|\nabla u|^2+\frac{1}{n-1}\int_{M}\varphi^2|\nabla|\nabla u||^2\leq\int_{M}|\nabla \varphi|^2|\nabla u|^2.
\end{align*}
for any $\varphi\in C^{1}_{0}(M)$, where $\Lambda_{n}=\frac{1}{1+\epsilon_{n}}-\frac{n-1}{n}\left(1+\frac{n-2}{\sqrt{n-1}}\frac{\epsilon_{n}}{1+\epsilon_{n}}\right)$.
\end{lemm}
\begin{proof}
Fix $p\in M$, choose an orthonormal basis $\{e_{i}\}_{i=1}^{n}$ in $T_{p}M$. Denote $A_{ij}=A(e_{i},e_{j})$. By the Gauss equation,
\begin{align*}
  \Ric_{M}(e_{1},e_{1})
  &=A_{11}H-A_{11}^2-\sum\limits_{i=2}^{n}A_{1i}^2.
\end{align*}
Using this, we estimate $|A|^2$ as follows:
\begin{align*}
  |A|^2 &\geq A_{11}^2+\sum\limits_{i=2}^{n}A_{ii}^2+2\sum\limits_{i=2}^{n}A_{1i}^2 \\
   & \geq A_{11}^2+\frac{1}{n-1}(\sum\limits_{i=2}^{n}A_{ii})^2+2\sum\limits_{i=2}^{n}A_{1i}^2\\
  &=A_{11}^2+\frac{1}{n-1}(H^2-2HA_{11}+A_{11}^2)+2\sum\limits_{i=2}^{n}A_{1i}^2\\
&\geq -\frac{n}{n-1}\Ric_{M}(e_{1},e_{1})+\frac{n-2}{n-1}HA_{11}\\
  &\geq -\frac{n}{n-1}\Ric_{M}(e_{1},e_{1})- \frac{n-2}{n-1}\delta_{n}|A|^2,
  \end{align*}
where  we used the fact $|H|\geq-\delta_{n}|A|$ in the last inequality.
Thus, we obtain a lower bound on the  Ricci curvature:
\begin{align}\label{Ric-inq}
\Ric_{M}(e_{1},e_{1})\geq-\frac{n-1}{n}\left(1+ \frac{n-2}{n-1}\delta_{n}\right)|A|^2.
\end{align}
Since $M^n$ is stable $F$-minimal, the pinching condition (\ref{cond-1}) implies that
$$\int_{M}\frac{1}{1+\epsilon_{n}}|A|^2f^2\leq\int_{M}|\nabla f|^2, \forall f\in C^{1}_{0}(M).$$
Replacing $f$ by $|\nabla u|\varphi$, we rewrite the $F$-stability inequality as below:
\begin{align}\label{eq-x1}
  \frac{1}{1+\epsilon_{n}}\int_{M}\varphi^2|\nabla u|^2|A|^2 & \leq\int_{M}|\nabla\varphi|^2|\nabla u|^2+2\int_{M}(\varphi|\nabla u|\bangle{\nabla\varphi,\nabla|\nabla u|}+\varphi^2|\nabla|\nabla u||^2) \\
  &=\int_{M}|\nabla\varphi|^2|\nabla u|^2-\int_{M}\varphi^2|\nabla u|\Delta|\nabla u|.\nonumber
\end{align}
On the one hand, by the improved Kato's inequality for harmonic functions, $$|\nabla^{2}u|^2\geq\frac{n}{4(n-1)}|\nabla u|^{-2}|\nabla|\nabla u|^2|^{2}=\frac{n}{n-1}|\nabla|\nabla u||^2.$$
Combined with the Bochner formula
\begin{align*}
\Delta|\nabla u|^2&=2\Ric_{M}(\nabla u,\nabla u)+2|\nabla^2 u|^2,
\end{align*} the improved Kato inequality for harmonic functions
$$|\nabla^{2}u|^2\geq\frac{n}{4(n-1)}|\nabla u|^{-2}|\nabla|\nabla u|^2|^{2}=\frac{n}{n-1}|\nabla|\nabla u||^2,$$ and the inequality (\ref{Ric-inq}) with $e_1=\frac{\nabla u}{|\nabla u|}$, we have
$$\Delta|\nabla u|\geq-\frac{n-1}{n}\left(1+ \frac{n-2}{n-1}\delta_{n}\right)|A|^{2}|\nabla u|+\frac{1}{n-1}|\nabla u|^{-1}|\nabla|\nabla u||^2.$$
Inserting into \eqref{eq-x1}, we get the assertion.
\end{proof}
\begin{rema}\label{rema-2}
$\Lambda_{n}>0$ if $\epsilon_n>0$ is sufficiently small. In particular, in dimensions  $n=4,5$, we can choose the following pinching constant:
\begin{itemize}
  \item $n=4$, let $1+\epsilon_{4}=\frac{23}{20}$, then $\Lambda_{4}\approx 0.0335$. 
  \item $n=5$, let $1+\epsilon_{5}=\frac{1001}{1000}$, then $\Lambda_{5}\approx 0.1998$.
\end{itemize}
\end{rema}
By using a standard argument as Cao-Shen-Zhu \cite{Cao-Shen-Zhu1997}, we get the one-endness of $M$.
\begin{prop}\label{one-end}
  Assume $F$ satisfies (\ref{cond-1}) such that $\Lambda_n>0$, where $\Lambda_n$ is given in Lemma \ref{one-end-improve}. Then any complete, two-sided stable $F$-minimal immersion $M^{n}$ in $\bR^{n+1}$ has only one end.
\end{prop}
\begin{proof}
  We argue by contradiction. If not, then $M^{n}$ has at least two ends. Therefore, Proposition \ref{least-two ends} implies that $M^{n}$ admits a nontrivial harmonic function $u$ with finite Dirichlet energy, i.e. $\int_{M}|\nabla u|^2\leq C<\infty$. For $R>0$, take $\varphi\in C^{1}_{0}(M)$, such that $\varphi|_{B_{R}(0)}=1$, $\varphi|_{B_{2R}\setminus B_{R}}=0$ and $|\nabla \varphi|\leq\frac{2}{R}$. Then Lemma \ref{one-end-improve} implies that
  \[
  \int_{B_{R}}\Lambda_n|A|^2|\nabla u|^2+\frac{1}{n-1}|\nabla|\nabla u||^2\leq\frac{4}{R^2}\int_{M}|\nabla u|^2\leq\frac{4C}{R^2}.
  \]
  Sending $R\to\infty$, we conclude that $|\nabla|\nabla u||^2=0.$
  In particular, this implies that $|\nabla u|$ is a constant. But $u$ is not a constant, which means that $|\nabla u|>0$. Hence, this implies that $$\int_{M}1=\frac{1}{|\nabla u|^2}\int_{M}|\nabla u|^2<\infty,$$
  contradicts with Proposition \ref{infinite volume}.
\end{proof}

\section{A conformal deformation of metrics}
Let $M^n\to\bR^{n+1} (n=4,5)$ be a complete two-sided stable $F$-minimal immersed hypersurface passing through $0\in \bR^{n+1}$.
Let $g$ be the induced metric on $M$ and consider the conformal deformation by Gulliver-Lawson \cite{GL} on $M\setminus\{0\}$, i.e., $N=M\setminus\{0\}$ with metric $\tilde{g}=r^{-2}g$, where $r(x)=dist_{\bR^{n+1}}(0, x)$. We denote the related notations in $(N, \tilde{g})$ by a tilde.
We recall that the weighted bi-Ricci curvature $\biRic_{\alpha}$ is defined as:
$$\biRic_{\alpha}(e_{1},e_{2})=\sum\limits_{i=2}^{n}\Rm(e_{1},e_{i},e_{i},e_{1})+\alpha\sum\limits_{j=3}^{n}\Rm(e_{2},e_{j},e_{j},e_{2}),$$
and the minimum of $\biRic_{\alpha}$ is denoted by $$\lambda_{\biRic_{\alpha}}=\min_{v, w\in T_{p}M,~v\perp w,~|v|=|w|=1}\biRic_{\alpha}(v,w).$$
In the case $\alpha=1$, $\biRic_{\alpha}$ is the bi-Ricci curvature, which was first introduced by Shen-Ye in \cite{Shen-Ye96}.

The aim of this section is to prove the following
\begin{prop}\label{low-bound}
Let $n=4, 5$. Assume $F$ satisfies (\ref{cond-1}). Let $M^n\to\bR^{n+1}$ be a complete two-sided stable $F$-minimal immersed hypersurface passing through $0\in \bR^{n+1}$. Then for any $\varphi\in C^{1}_{0}(N,\tilde{g})$,
\begin{eqnarray}\label{ineq-x1}\int_{N}|\tilde{\nabla} \varphi|^2_{\tilde{g}}d\tilde{\mu}\geq \int_{N}(\tau_{n}-\eta_{n}\tilde{\lambda}_{\biRic_{\alpha}})\varphi^2d\tilde{\mu},\end{eqnarray}
  where $\tau_n>0$ and $\eta_n>0$ are given by
  \begin{itemize}
  \item $n=4$, $\quad \tau_4=\frac{7}{4}\cdot\frac{1+\epsilon_{4}-6\epsilon_{4}^2}{d_{4}(1+\epsilon_{4})^2}-\frac{9}{8}, \quad \eta_4=\frac{1+\epsilon_{4}-6\epsilon_{4}^2}{d_{4}(1+\epsilon_{4})^2},\quad d_{4}\approx 1.051.$
  \item $n=5$, $\quad \tau_5=\frac{5183}{1600}\cdot\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2}-\frac{39}{16}, \quad \eta_5=\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2},\quad d_{5}\approx 1.10536028.$
\end{itemize}
\end{prop}

Let $\overrightarrow{H}=-H\nu$, $\overrightarrow{x}$ denotes the position vector field in $\bR^{n+1}$. 
We first recall several formulas under the conformal deformation, which have been given in \cite{Chodosh-Li-Minter-Stryker-5bernstein, Mazet-6Bernstein}.
 \begin{prop}\label{conf-indentity}
   Let $(M^n,g)$ be a complete two-sided, stable $F$-minimal hypersurface in $\bR^{n+1}$, then the following equalities hold:\\
$(i)$. $$\tilde{\Delta}(\log r)=n-n|dr|^2+\small\bangle{\overrightarrow{H},\overrightarrow{x}},$$
$(ii)$.
$$r^{2}\Rm(e_{i},e_{j},e_{j},e_{i})={\tilde{\Rm}}(e_{i},e_{j},e_{j},e_{i})-2+|dr|^2+(dr(e_{i}))^2+(dr(e_{j}))^2+\bangle{\overrightarrow{x},\nu}(A_{ii}+A_{jj}).$$
 \end{prop}
{Using the Gauss equation and Proposition \ref{conf-indentity} (ii), we have expressions for the weighted bi-Ricci curvatures under the two different metrics. }
\begin{prop}\label{pro-2}
   \begin{align*}
   {\biRic}_{\alpha}(e_{1},e_{2})&=\sum\limits_{i=2}^{n}(A_{ii}A_{11}-A_{i1}^2)+\alpha\sum\limits_{j=3}^{n}(A_{22}A_{jj}-A_{2j}^2).\\
   \widetilde{\biRic}_{\alpha}(\tilde{e}_{1},\tilde{e}_{2})
 &=r^2(\sum\limits_{i=2}^{n}(A_{11}A_{ii}-A_{1i}^2)+\alpha\sum\limits_{j=3}^{n}(A_{22}A_{jj}-A_{2j}^2))\\
 &\quad+2((n-1)+\alpha(n-2))-(n+\alpha(n-1))|dr|^2 \\
&\quad-((n-2-\alpha)dr(e_{1})^2+\alpha(n-3)dr(e_{2})^2)\\
&-\bangle{\overrightarrow{x},\nu}((n-2-\alpha)A_{11}+\alpha(n-3)A_{22}+(1+\alpha)H).
 \end{align*}
 \end{prop}

We select the basis vectors $\tilde{e}_1$ and $\tilde{e}_2$ such that $\widetilde{\biRic_{\alpha}}(\tilde{e}_{1},\tilde{e}_2)=\tilde{\lambda}_{\biRic_{\alpha}}$.
Assume $\alpha\leq 1$ and $a\in \bR$ satisfy $a\geq\frac1 2$ and $2a\geq\alpha$. Let $\varsigma_{n}$ be a positive constant, then
 \begin{align*}
&ar^2|A|^2+r^2(\sum\limits_{i=2}^{n}(A_{11}A_{ii}-A_{1i}^2)+\alpha\sum\limits_{j=3}^{n}(A_{22}A_{jj}-A_{2j}^2))\\
 &\quad-\bangle{\overrightarrow{x},\nu}((n-2-\alpha)A_{11}+\alpha(n-3)A_{22}+(1+\alpha)H)\\
\geq & a r^2\sum\limits_{i=1}^{n}A_{ii}^2+r^2(\sum\limits_{i=2}^{n}A_{11}A_{ii}+\alpha\sum\limits_{j=3}^{n}A_{22}A_{jj})\\
 &\quad-\bangle{\overrightarrow{x},\nu}((n-2-\alpha)A_{11}+\alpha(n-3)A_{22}+(1+\alpha)H)\\
 \geq & a r^2\sum\limits_{i=1}^{n}A_{ii}^2+r^2(\sum\limits_{i=2}^{n}A_{11}A_{ii}+\alpha\sum\limits_{j=3}^{n}A_{22}A_{jj})\\
 &\quad-\bangle{\overrightarrow{x},\nu}((n-2-\alpha)A_{11}+\alpha(n-3)A_{22}+H)-\varsigma_{n}\alpha^{2}r^{2}H^{2}-\frac{1}{4\varsigma_{n}}\\
     :=&q^{T}Bq+C^Tq-\varsigma_{n}\alpha^{2}r^{2}H^{2}-\frac{1}{4\varsigma_{n}},
 \end{align*}
where 
 \begin{equation*}
 B=
 \begin{pmatrix}
   a & \frac{1}{2} &  \frac{1}{2} &  \cdots &  \frac{1}{2} \\
    \frac{1}{2} & a & \frac{\alpha}{2} & \cdots & \frac{\alpha}{2} \\
    \frac{1}{2} & \frac{\alpha}{2} & a & \cdots & 0 \\
   \vdots & \vdots & \vdots & \ddots & \vdots \\
    \frac{1}{2} & \frac{\alpha}{2} & 0 & \cdots & a\\
 \end{pmatrix},
 \quad q=r
 \begin{pmatrix}
   A_{11} \\
   A_{22} \\
   A_{33}\\
   \vdots \\
   A_{nn}\\
 \end{pmatrix},
\end{equation*}
and
 \begin{equation*}
   \quad C=-\bangle{\frac{\overrightarrow{x}}{r},\nu}
   \begin{pmatrix}
     n-1-\alpha \\
     \alpha(n-3)+1 \\
    1 \\
    \vdots \\
    1
   \end{pmatrix}=-\bangle{\frac{\overrightarrow{x}}{r},\nu}\bar{C}.
 \end{equation*}
 If $B$ is positive definite, by using the inequality
  \begin{align*}
 q^{T}Bq+C^{T}q\geq -\frac{1}{4}C^{T}B^{-1}C,
 \end{align*} we have
 \begin{align*}
&\varsigma_{n}\alpha^{2}r^{2}H^{2}+ar^2|A|^2+r^2(\sum\limits_{i=2}^{n}(A_{11}A_{ii}-A_{1i}^2)+\alpha\sum\limits_{j=3}^{n}(A_{22}A_{jj}-A_{2j}^2))\\
 &\quad-\bangle{\overrightarrow{x},\nu}((n-2-\alpha)A_{11}+\alpha(n-3)A_{22}+H)-\frac{1}{4\varsigma_{n}}\\
&\geq-b\bangle{\frac{\overrightarrow{x}}{r},\nu}^2-\frac{1}{4\varsigma_{n}},
 \end{align*}
 where  $b:=\frac{1}{4}\bar{C}^{T}B^{-1}\bar{C}$.
 Therefore, since we choose $\alpha\le 1$,  $H^{2}\leq \delta_{n}^{2}|A|^{2}$, let $d_{n}=a+\varsigma_{n}\alpha^{2}\delta_{n}^{2}$, then
 \begin{align*}
   d_{n}r^2|A|^2+\widetilde{\biRic}_{\alpha}(\tilde{e}_{1},\tilde{e}_{2}) &\geq-b\bangle{\frac{\overrightarrow{x}}{r},\nu}^2+2((n-1)+\alpha(n-2))-(n+\alpha(n-1))|dr|^2 \\
&\quad-((n-2-\alpha)dr(e_{1})^2+\alpha(n-3)dr(e_{2})^2)-\frac{1}{4\varsigma_{n}}\\
&\geq-b\bangle{\frac{\overrightarrow{x}}{r},\nu}^2+2((n-1)+\alpha(n-2))-(n+\alpha(n-1))|dr|^2 \\
&\quad-(n-2-\alpha)|dr|^2+(n-2-\alpha(n-2))dr(e_{2})^2-\frac{1}{4\varsigma_{n}}\\
&\geq-b(1-|dr|^2)+2(n-1+\alpha(n-2))-(2n-2+\alpha(n-2))|dr|^2-\frac{1}{4\varsigma_{n}}.
 \end{align*}
We reformulate the expression as follows:
\begin{prop}\label{pro-3}
  Let $n=4, 5$, and $a\geq\frac{1}{2}$, $2a\geq\alpha$, $\alpha\leq 1$, $\varsigma_{n}>0$. If $B$ is a positive definite matrix, then the following inequality holds:
  $$r^2|A|^2\geq-\frac{b}{d_{n}}(1-|dr|^2)+\frac{2}{d_{n}}(n-1+\alpha(n-2))-\frac{1}{d_{n}}(2n-2+\alpha(n-2))|dr|^2-\frac{1}{d_{n}}\tilde{\lambda}_{\biRic_\alpha}-\frac{1}{4\varsigma_{n}d_{n}}.$$
\end{prop}
We now adapt the estimate for $|A|^2$ from Proposition \ref{pro-3} to give a characterization of $F$-stability in terms of the $\alpha$-bi-Ricci curvature under the conformal deformation.

\begin{proof}[Proof of Proposition \ref{low-bound}]
On the one hand, under the conformal metric $\tilde{g}$, we have
  $$d\tilde{\mu}=r^{-n}d\mu \quad and \quad |\tilde{\nabla} f|^2_{\tilde{g}}=r^2|\nabla f|^2.$$
The $F$-stability inequality yields
   $$\int_{N}r^{n-2}|\tilde{\nabla} f|^2_{\tilde{g}}d\tilde{\mu}\geq\frac{1}{1+\epsilon_{n}}\int_{N}r^{n-2}(r^2|A|^2)f^2d\tilde{\mu},$$\
   for any $f\in C^{1}_{0}(N,\tilde{g})$. Take $f=r^{\frac{2-n}{2}}\varphi$ for $\varphi\in C^{1}_{0}(N,\tilde{g})$, then
   $$\tilde{\nabla} f=r^{\frac{2-n}{2}}\tilde{\nabla}\varphi-\frac{n-2}{2}r^{-\frac{n}{2}}\varphi\tilde{\nabla}r,$$
and
   $$|\tilde{\nabla f}|^2_{\tilde{g}}=r^{2-n}|\tilde{\nabla} \varphi|^2_{\tilde{g}}+\frac{(n-2)^2}{4}r^{-n}\varphi^2|\tilde{\nabla} r|^2_{\tilde{g}}-(n-2)r^{1-n}\varphi\bangle{\tilde{\nabla}\varphi,\tilde{\nabla}r}_{\tilde{g}}=:L_{1}+L_{2}+L_{3}.$$
   We obtain $$\int_{N}r^{n-2}L_{1}d\tilde{\mu}=\int_{N}|\tilde{\nabla}\varphi|^2_{\tilde{g}}d\tilde{\mu},$$
   and
   $$\int_{N}r^{n-2}L_{2}d\tilde{\mu}=\int_{N}\frac{(n-2)^2}{4}|dr|^2\varphi^2d\tilde{\mu},$$
    where we used the fact that $r^{-2}|\tilde{\nabla} r|^2_{\tilde{g}}=|dr|^2$.
   Integrating by parts, the third term yields
   \begin{align*}
     \int_{N}r^{n-2}L_{3}d\tilde{\mu}&=-\int_{N}\frac{n-2}{2}\bangle{\tilde{\nabla}(\varphi^2),\tilde{\nabla}(\log r)}_{\tilde{g}}d\tilde{\mu}\\
     &=\int_{N}\frac{n-2}{2}\tilde{\Delta}(\log r)\varphi^2d\tilde{\mu}\\
     &=\int_{N}(\frac{n(n-2)}{2}-\frac{n(n-2)}{2}|dr|^2+\frac{n-2}{2}\bangle{\overrightarrow{H},\overrightarrow{x}})\varphi^2d\tilde{\mu}.
    \end{align*}
    Put them together, by Lemma \ref{mean-estimate}, we obtain
    \begin{align*}
      \int_{N}|\tilde{\nabla} \varphi|^2_{\tilde{g}}d\tilde{\mu}&\geq \int_{N}\bigg(\frac{1}{1+\epsilon_{n}} r^2|A|^2-\frac{n(n-2)}{2}+\frac{(n-2)(n+2)}{4}|dr|^2-\frac{n-2}{2}\bangle{\overrightarrow{H},\overrightarrow{x}}\bigg)\varphi^2d\tilde{\mu}\\
      &\geq\int_{N}\bigg(\frac{1}{1+\epsilon_{n}} r^2|A|^2-\frac{n(n-2)}{2}+\frac{(n-2)(n+2)}{4}|dr|^2-(n-2)H^2r^2-\frac{n-2}{16}\bigg)\varphi^2d\tilde{\mu}\\
      &\geq\int_{N}\bigg(\frac{(1+\epsilon_{n})-(n-2)(n-1)\epsilon_{n}^2}{(1+\epsilon_{n})^2}r^2|A|^2+\frac{(n-2)(n+2)}{4}|dr|^2-\frac{(8n+1)(n-2)}{16}\bigg)\varphi^2d\tilde{\mu}.
    \end{align*}

 On the other hand, combining with Proposition \ref{pro-3}, we consider the following two cases:
 
\textbf{Case 1.} For $n=4$, $1+\epsilon_{4}=\frac{23}{20}$, choosing $a=\alpha=1$, $\varsigma_{4}=1$, by software Mathematica, we can easily check $B$ is positive definite and $b=\frac{3}{2}$, $d_{4}\approx1.051$, then the above inequality yields:
\begin{align*}
&\frac{1+\epsilon_{4}-6\epsilon_{4}^2}{(1+\epsilon_{4})^2}r^2|A|^2-\frac{33}{8}+3|dr|^2\\
&\geq \frac{1+\epsilon_{4}-6\epsilon_{4}^2}{d_{4}(1+\epsilon_{4})^2}\bigg(\frac{33}{4}-\frac{13}{2}|dr|^2-\tilde{\lambda}_{\biRic_{\alpha}}\bigg)-\frac{33}{8}+3|dr|^2\\
&= \frac{33}{4}\cdot\frac{1+\epsilon_{4}-6\epsilon_{4}^2}{d_{4}(1+\epsilon_{4})^2}-\frac{33}{8}+\big(3-\frac{13}{2}\cdot\frac{1+\epsilon_{4}-6\epsilon_{4}^2}{d_{4}(1+\epsilon_{4})^2}\big)|dr|^2-\frac{1+\epsilon_{4}-6\epsilon_{4}^2}{d_{4}(1+\epsilon_{4})^2}\tilde{\lambda}_{\biRic_\alpha}.\\
 &\geq \frac{7}{4}\cdot\frac{1+\epsilon_{4}-6\epsilon_{4}^2}{d_{4}(1+\epsilon_{4})^2}-\frac{9}{8}-\frac{1+\epsilon_{4}-6\epsilon_{4}^2}{d_{4}(1+\epsilon_{4})^2}\tilde{\lambda}_{\biRic_\alpha}\\
 &:=\tau_{4}-\eta_{4}\tilde{\lambda}_{\biRic_\alpha},
\end{align*}
where $\tau_{4}\approx 0.1529$, $\eta_{4}\approx 0.7303$. In the second inequality, we used the fact that $$3-6.5\cdot\frac{1+\epsilon_{4}-6\epsilon_{4}^2}{d_{4}(1+\epsilon_{4})^2}\approx-1.7466<0$$ and
$|dr|\le 1$. 

\textbf{Case 2.} For $n=5$, $1+\epsilon_{5}=\frac{1001}{1000}$, 
if we choose $a=\frac{217}{200}$, $\alpha=\frac{11}{12}$, $\varsigma_{5}=160$, then $b=5.0311$ and $B$ is positive definite, $d_{5}\approx 1.0855$, the above inequality leads to:
\begin{align*}
&\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{(1+\epsilon_{5})^2}r^2|A|^2-\frac{123}{16}+\frac{21}{4}|dr|^2\\
&\geq\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2}(-b+8+6\alpha-\frac{1}{64})-\frac{123}{16}+\bigg(\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2}(b-(8+3\alpha))+\frac{21}{4}\bigg)|dr|^2\\
 &\quad-\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2}\tilde{\lambda}_{\biRic_{\alpha}}\\
&= (\frac{651}{200}-\frac{1}{64})\cdot\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2}-\frac{39}{16}-\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2}\tilde{\lambda}_{\biRic_{\alpha}}\\
&\geq \frac{5183}{1600}\cdot\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2}-\frac{39}{16}-\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2}\tilde{\lambda}_{\biRic_{\alpha}}\\
 &:= \tau_{5}-\eta_{5}\widetilde{\lambda}_{\biRic_{\alpha}},
 \end{align*}
 where $\tau_{5}\approx 0.09181$, $\eta_{5}\approx 0.92027$.  In the second inequality, we used the fact that $$\frac{1+\epsilon_{5}-12\epsilon_{5}^2}{d_{5}(1+\epsilon_{5})^2}(b-(8+3\alpha))+\frac{21}{4}\approx-0.012942371<0$$ and
$|dr|\le 1$. 
\end{proof}

\section{$\mu$-bubbles and proof of Theorem 2}
In this section, under the condition of spectral uniformly positivity of the $\alpha$-bi-Ricci curvature on $(N,\tilde{g})$, we construct a warped $\mu$-bubble whose volume has an upper bound. We follow closely Chodosh-Li-Minter-Stryker \cite{Chodosh-Li-Minter-Stryker-5bernstein} and Mazet \cite{Mazet-6Bernstein}. 

First, by \cite[Theorem 1]{FCS80}, there exists a positive function $w$ on $(N,\tilde{g})$ satisfying
\begin{align}\label{JACfun}
-\tilde{\Delta}w=(\tau_{n}-\eta_{n}\tilde{\lambda}_{\biRic_{\alpha}})w.
\end{align}
We use $w$ for the weight function in $\cA_{k}(\Omega)$ as below.

\subsection{Warped $\mu$-bubbles}
Given an $n$-dimensional compact Riemannian manifold $(M_{0},\tilde{g})$ with boundary consisting of two disjoint components, denoted by $\partial M_{0}=\partial_{+}M_{0}\cup\partial_{-}M_{0}$, where neither of them are empty. Let $h$ be a smooth function on $\mathring{M_{0}}$(interior of $M_0$) with $h\to \pm\infty$ on $\partial_{\pm}M_{0}$, choose a finite perimeter set $\Omega_{0}$ with $\partial\Omega_{0}\subset\mathring{M_{0}}$ and $\partial_{+}M_{0}\subset\Omega_{0}$. We consider the following warped prescribed mean curvature functional $\cA_{k}$ defined by $$\cA_{k}({\Omega})=\int_{\partial^{*} \Omega}w^{k}d\tilde{\mu}^{n-1}-\int_{M_{0}}(\chi_{\Omega}-\chi_{\Omega_{0}})hw^{k}d\tilde{\mu}^{n}.$$

If there exists a finite perimeter set $\Omega$ minimizing $\cA_{k}$ among the Caccioppoli sets containing $\partial_{+}M_{0}$, we call $\Sigma=\partial^{*}\Omega$ a warped $\mu$-bubble, where $\partial^{*}\Omega$ denotes the reduced boundary of $\Omega$.  Let $\tilde{\nu}$ be the outward unit normal vector field of $\Sigma$. We note that the $\mu$-bubble was first introduced by Gromov \cite{Gromov2019Fourlecture}, whose existence and regularity theory have been established in  \cite{general soap bubbles,zhujintian2021}.
The first and second variation formulas are as follows. We refer to \cite{Hong-cmc nonexis,Mazet-6Bernstein} for details.
\begin{prop}[first and second variation formula]
  Let $\Omega_{t}$ be a smooth family of sets of finite perimeter with $\Omega_{t}\big|_{t=0}=\Omega$ and with velocity $\psi\tilde{\nu}$ at $t=0$, then $$\frac{d}{dt}\big|_{t=0}\cA_{k}(\Omega_{t})=\int_{\Sigma}(\tilde{H}_\Sigma w^{k}+\bangle{\tilde{\nabla}_{M_0} w^{k},\tilde{\nu}}-hw^{k})\psi d\tilde{\mu}^{n-1},$$
  where $\tilde{H}_\Sigma$ is the mean curvature of $\Sigma$. In particular, the $\mu$-bubble $\Sigma$ satisfies $$\tilde{H}_\Sigma=-kw^{-1}\bangle{\tilde{\nabla}_{M_0}w,\tilde{\nu}}+h.$$Also,
\begin{align*}
  \frac{d^2}{dt^2}\bigg|_{t=0}\cA_{k}(\Omega)
&=\int_{\Sigma}w^{k}\bigg(|\tilde{\nabla}_{\Sigma}\psi|^2-(|\tilde{A}_{\Sigma}|^2+\tilde{\Ric}_{M_{0}}(\tilde{\nu},\tilde{\nu}))\psi^2-kw^{-2}\bangle{\tilde{\nabla}_{M_{0}} w,\tilde{\nu}}^2\psi^2\\
  &\quad+kw^{-1}(\tilde{\Delta}_{M_{0}}w-\tilde{\Delta}_{\Sigma}w-\tilde{H}_\Sigma\bangle{\tilde{\nabla}_{M_0} w,\tilde{\nu}})\psi^2+\bangle{\tilde{\nabla}_{M_0} h,\tilde{\nu}}\psi^2\bigg)d\tilde{\mu}^{n-1}.
  \end{align*}
\end{prop}

\subsection{Volume estimates of $\mu$-bubbles}
For manifolds satisfying a spectral Ricci curvature lower bound, Chodosh-Li-Minter-Stryker \cite{Chodosh-Li-Minter-Stryker-5bernstein} and Antonelli-Xu in \cite{Antonelli-Xu} established the following spectral Bishop-Gromov volume comparison theorem.
\begin{theo}[\cite{Chodosh-Li-Minter-Stryker-5bernstein, Antonelli-Xu}]\label{Antonelli-Xu}
 Let $\Sigma^n$, $n\geq 3$, be an $n$-dimensional simply connected compact smooth manifold, and let $0\leq \theta\leq \frac{n-1}{n-2}$, $\gamma>0$. Assume there is a positive function $\omega\in C^{\infty}(\Sigma)$ such that:
  $$\theta\Delta_{\Sigma}\omega\leq \lambda_{\Ric} \omega-(n-1)\gamma \omega,$$
where $\lambda_{\Ric}$ is the smallest eigenvalue of the  Ricci tensor on $\Sigma$, then there exists a sharp volume bound $$\Vol(\Sigma^{n})\leq\gamma^{-\frac{n}{2}}\Vol(\bS^{n}).$$
\end{theo}

In the following, we find good $\mu$-bubble in $(N, \tilde{g})$ and 
apply the spectral Bishop-Gromov volume comparison theorem to get the volume estimate for the $\mu$-bubble.
\begin{prop}\label{diam-volume bound}
  Let $n=4, 5$ and $(N_{0},\tilde{g})$ be a compact subset of $(N^{n},\tilde{g})$ with connected boundary. Suppose there exists a point $p\in N_{0}$ such that $d_{\tilde{g}}(p,\partial N_{0})\geq 10\pi$. Let $w$ be a positive function on $(N_{0},\tilde{g})$ such that
\begin{align}\label{con-10}
  -\tilde{\Delta}_N w\geq(\tau_{n}-\eta_{n}\tilde{\lambda}_{\biRic_{\alpha}})w.
\end{align}
  Then there exists a connected relative open set $\Omega$ containing $\partial N_{0}$ such that $\Omega\subset B^{\tilde{g}}_{10\pi}(\partial N_{0})$ and 
 $$\Vol_{\tilde{g}}(\Sigma)\leq\left(\frac{\tau_{n}}{2(n-2)\alpha_{n}\eta_{n}}\right)^{-\frac{n-1}{2}}\Vol(\bS^{n-1}),$$
  where $\Sigma$ is a connected component of $\partial\Omega\setminus\partial N_{0}$, $B^{\tilde{g}}_{10\pi}(\partial N_0)$ denotes the $10\pi$-neighborhood of $\partial N_0$ with respect to the metric $\tilde{g}$. Here $\alpha_{4}=1$, $\alpha_{5}=\frac{11}{12}$.
\end{prop}

\begin{proof}
According to the second variation formula, following the same computations from \cite[Section 4.2]{Mazet-6Bernstein}, and using the spectral $\alpha$-bi-Ricci condition (\ref{con-10}), we obtain
\begin{align*}
  \frac{4}{4-k}\int_{\Sigma}|\tilde{\nabla}_{\Sigma}\psi|^2 
  &\geq\int_{\Sigma}\psi^2\bigg(k\tau_{n}-\alpha_{n}\tilde{\lambda}_{\Ric}+|\tilde{A}_{\Sigma}|^2+\alpha_{n} \tilde{H}_{\Sigma}\tilde{A}_{11}^{\Sigma}-\alpha_{n}\sum\limits_{i=1}^{n-1}(\tilde{A}_{1i}^{\Sigma})^2+k(\tilde{d}\ln w(\tilde{\nu}))^2\\
  &\quad+ k\tilde{H}_{\Sigma}\tilde{d}\ln w(\tilde{\nu})+\tilde{d}h(\tilde{\nu})\bigg).
\end{align*}
Choose $k=\frac{1}{\eta_{n}}$, and set
\begin{align*}
L_{n}&=
|\tilde{A}_{\Sigma}|^2+\alpha_{n} \tilde{H}_{\Sigma}\tilde{A}_{11}^{\Sigma}-\alpha_{n}\sum\limits_{i=1}^{n-1}(\tilde{A}_{1i}^{\Sigma})^2+\frac{1}{k}(\tilde{H}_{\Sigma}-h)^2+\tilde{H}_{\Sigma}(h-\tilde{H}_{\Sigma}).
\end{align*}
If we can show that 
\begin{align}\label{Ln-estimate}
L_{n}>\beta_{n} h^2,
\end{align}
then
$$\frac{4}{4-k}\int_{\Sigma}|\tilde{\nabla}_{\Sigma}\psi|^2\geq\int_{\Sigma}\psi^2\left(\frac{\tau_{n}}{2\eta_{n}}-\alpha_{n}\tilde{\lambda}_{\Ric}\right)+\psi^2\left(\frac{\tau_{n}}{2\eta_{n}}+\beta_{n} h^2+\tilde{d}h(\tilde{\nu})\right).$$
Following the method of \cite{general soap bubbles,Chodosh-Li-Minter-Stryker-5bernstein,Hong-cmc nonexis,Mazet-6Bernstein}, we can construct a smooth function $h$ such that $\frac{\tau_{n}}{2\eta_{n}}+\beta_{n} h^2-|\tilde{\nabla}_{N}h|>0$ holds on $\Sigma$.
Therefore, there exists a positive Jacobi function $\omega$, such that $$\frac{4\eta_{n}}{(4\eta_{n}-1)\alpha_{n}}\tilde{\Delta}_{\Sigma}\omega\leq\left(\tilde{\lambda}_{\Ric}-\frac{\tau_{n}}{2\alpha_{n}\eta_{n}}\right)\omega.$$
To prove inequality (\ref{Ln-estimate}) holds, we argue as in \cite[Section 4.2]{Mazet-6Bernstein}, write $L_{n}$ as a quadratic form and denoted $S_{n}$ by the associated matrix. Then, it suffices to show
\begin{equation*}
  S_{n}=
  \begin{pmatrix}
    \frac{1}{n-1}+\frac{\alpha_{n}}{n-1}-\frac{\alpha_{n}}{(n-1)^2}+\frac{1}{k}-1 & \frac{\alpha_{n}}{2}\sqrt{\frac{n-2}{n-1}}(1-\frac{2}{n-1}) & \frac{1}{2}-\frac{1}{k} \\
    \frac{\alpha_{n}}{2}\sqrt{\frac{n-2}{n-1}}(1-\frac{2}{n-1}) & 1-\frac{n-2}{n-1}\alpha_{n} & 0\\
    \frac{1}{2}-\frac{1}{k} & 0 & \frac{1}{k}-\beta_{n}\\
  \end{pmatrix}
  >0.
\end{equation*}
We again divide the discussion into two cases.

\textbf{Case 1.}
For $n=4$, the matrix $S_{4}$ becomes
\begin{equation*}
  S_{4}=
  \begin{pmatrix}
    \frac{2\alpha_{4}}{9}+\frac{1}{k}-\frac{2}{3} & \frac{\alpha_{4}}{6}\sqrt{\frac{2}{3}} & \frac{1}{2}-\frac{1}{k} \\
    \frac{\alpha_{4}}{6}\sqrt{\frac{2}{3}} & 1-\frac{2}{3}\alpha_{4} & 0\\
    \frac{1}{2}-\frac{1}{k} & 0 & \frac{1}{k}-\beta_{4}\\
  \end{pmatrix}.
\end{equation*}
According to Proposition \ref{low-bound}, $\alpha_{4}=1$, $\tau_{4}\approx 0.1529$, let $\frac{1}{k}=\eta_{4}\approx0.7303$. If we choose $\beta_{4}=\frac{2}{5}$, using software Mathematica, we find $S_{4}$ is positive definite, then $L_{4}>\frac{2}{5}h^2$.

 Since  $$\frac{4\eta_{4}}{(4\eta_{4}-1)\alpha_{4}}\approx1.5205\leq\frac{n-2}{n-3}=2,$$
we can use Theorem \ref{Antonelli-Xu} to conclude  that 
$$\Vol_{\tilde{g}}(\Sigma^3)\leq (\frac{\tau_{4}}{4\alpha_{4}\eta_{4}})^{-\frac{3}{2}}\Vol(\bS^3)\leq 168\pi^2.$$

\textbf{Case 2.} For $n=5$, the matrix $S_{5}$ yields
\begin{equation*}
  S_{5}=
  \begin{pmatrix}
    \frac{3\alpha_{5}}{16}+\frac{1}{k}-\frac{3}{4} & \frac{\alpha_{5}}{4}\sqrt{\frac{3}{4}} & \frac{1}{2}-\frac{1}{k} \\
    \frac{\alpha_{5}}{4}\sqrt{\frac{3}{4}} & 1-\frac{3}{4}\alpha_{5} & 0\\
    \frac{1}{2}-\frac{1}{k} & 0 & \frac{1}{k}-\beta_{5}\\
  \end{pmatrix}.
\end{equation*}
Note that $\alpha_{5}=\frac{11}{12}$, $\tau_{5}\approx 0.09181$, choose $\frac{1}{k}=\eta_{5}\approx 0.92027$, and let $\beta_{5}=\frac{1}{10}$, we conclude that $S_{5}$ is positive definite, then $L_{5}>\frac{1}{10}h^2$.

Similarly, since $$\frac{4\eta_{5}}{(4\eta_{5}-1)\alpha_{5}}\approx 1.497799542\leq \frac{n-2}{n-3}=\frac{3}{2},$$ we can use Theorem \ref{Antonelli-Xu} to conclude  that $$\Vol_{\tilde{g}}({\Sigma}^4)\leq(\frac{\tau_{5}}{6\alpha_{5}\eta_{5}})^{-2}\Vol(\bS^{4})\leq 8200\pi^2.$$
\end{proof}

\subsection{Volume growth of the geodesic ball}
In this subsection, we will complete the proof of Theorem \ref{main-theo}. Since all preceding results are obtained under the conformal metric $\tilde{g}$, we now estimate the volume of $B_{R}(p)$. 

\begin{proof}[Proof of Theorem \ref{main-theo}]
We first assume that $M$ is simply connected.
Up to a translation, we shall identify $p$ with the origin $0$, and let $\bar{r}(x)=\dis_{M,g}(x,0)$. For a fixed sufficiently large $R>0$, we consider the geodesic ball $B_{e^{10\pi}R}(0)$ with respect to the metric $g$. By Proposition \ref{one-end}, $M\setminus B_{e^{10\pi}R}(0)$ must have only one unbounded connected component $E$. We set $M^{\prime}=M\setminus E$. As argued in the proof of \cite[Theorem 1.4]{Chodosh-Li-anisotropic}, by using simply-connectedness of $M$, we see $\partial M^{\prime}=\partial E$ is connected.

Applying Proposition \ref{diam-volume bound} to $(M^{\prime}\setminus\{0\},\tilde{g})$, there exists a connected open set $\Omega$ in the $10\pi$-neighborhood of $\partial M^{\prime}$, such that each connected component of $\partial \Omega\setminus\partial M^{\prime}$ has volume bounded by $(\frac{\tau_{n}}{2(n-2)\alpha_{n}\eta_{n}})^{-\frac{n-1}{2}}\Vol(\bS^{n-1})$.

Let $M_{1}$ be the connected component of $M^{\prime}\setminus\Omega$ that contains $0$. Analogously, we can apply the same argument as in \cite{Chodosh-Li-anisotropic} to show that $M\setminus M_{1}$ is connected, and $M_{1}$ has only one boundary component denoted by $\Sigma=\partial M_{1}$. By the second item in \cite[Lemma 6.2]{Chodosh-Li-anisotropic}, we have $\min\limits_{x\in\Sigma}\bar{r}(x)\geq R$, so $B_{R}(0)\subset M_{1}$. On the other hand, by comparing the intrinsic distance $\bar{r}$ with the extrinsic distance $r$, we have $\max\limits_{x\in\Sigma}r(x)\leq e^{10\pi} R$, so 
\begin{align*}\Vol_{g}(\Sigma)&=\int_{\Sigma}r^{n-1} d\tilde{\mu}^{n-1}\leq e^{10(n-1)\pi}R^{n-1}\Vol_{\tilde{g}}(\Sigma)\\&\leq e^{10(n-1)\pi}\left(\frac{\tau_{n}}{2(n-2)\alpha_{n}\eta_{n}}\right)^{-\frac{n-1}{2}}\Vol(\bS^{n-1})R^{n-1}.\end{align*}
Finally, Proposition \ref{iso-perimetric} implies that \begin{align*}\Vol_{g}(B_{R}(0))&\leq  \Vol_{g}(M_{1})\leq\frac{\|F\|_{C^{1}(\bS^{n})}}{n\min\limits_{\nu\in\bS^{n}}F(\nu)}e^{10\pi}R \Vol_{g}(\Sigma)\\&\leq \frac{e^{10n\pi}\left(\frac{\tau_{n}}{2(n-2)\alpha_{n}\eta_{n}}\right)^{-\frac{n-1}{2}}\Vol(\bS^{n-1})\|F\|_{C^{1}(\bS^{n})}}{n\min\limits_{\nu\in\bS^{n}}F(\nu)}R^n.\end{align*}
\begin{itemize}
    \item $n=4$,\quad\quad $C(F)=\frac{42\pi^2e^{40\pi}\|F\|_{C^{1}(\bS^{4})}}{\min\limits_{\nu\in\bS^{4}}F(\nu)}.$ 
    \item  $n=5$, \quad\quad$C(F)=\frac{1640\pi^2e^{50\pi}\|F\|_{C^{1}(\bS^{5})}}{\min\limits_{\nu\in\bS^{5}}F(\nu)}$. 
\end{itemize}
 This completes the proof in the case that $M$ is simply connected.

When $M$ is not simply connected, we lift it to its universal cover $\tilde{M}\subset \mathbb{R}^{n+1}$. We remark that in general $F$-stability may not hold on $\tilde{M}$. Nevertheless, by the local isometry of $M$ and $\tilde{M}$, we see that the local property \eqref{Hsmall} holds on $\tilde{M}$. Also by the argument of Fischer Colbrie-Schoen \cite{FCS80}, the modified stability inequality \eqref{delta-stable} holds on $\tilde{M}$. By using \eqref{Hsmall} and \eqref{delta-stable}, we can use the same argument to show that $\tilde{M}$ has only one end and it holds \eqref{ineq-x1} on $\tilde{M}$. This enables to apply Proposition \ref{diam-volume bound} and the rest of the proof is the same. The proof is completed.
\end{proof}

\

\begin{Ackn}
{\rm The authors would like to express their deep gratitude to the referee for his/her carefully reading and valuable suggestions which improve the context of this paper.}
\end{Ackn}

\

\bibliographystyle{plain}

\end{document}